calculate the weight of an object of mass 10kg on earth. what will be its mass and weight on the moon? (given g on earth is 10m/s^2

Grade:10

1 Answers

Aarti Gupta
9 years ago
Given-
Mass of object on earth = 10kg
Acceleration due to gravity,g = 10m/s2
Weight of object on earth = ?
Weight of object on moon = ?
Mass of object on moon = ?
Solution -
We know that weight of an object is given as,
w = mg
thus, putting the values,
w = 10 * 10 = 100kg
so,weight of object on earth = 100kg
Now,in order to calculate weight on moon,g on moon is 1/6 of g on earth
so,
w = m * g*1/6
w = 10 *10*1/6 = 100/6 = 16.666,or 16.67kg
Thus weight on moon = 16.67kg
Mass of an object does not changes when it is takes to moon ,it will remain same as it was on earth.
So,mass will be 10kg on moon also.